OpenStudy (anonymous):

suppose that a rumor is spreading in the united states that the airlines will soon be giving away free promotional tickets for us flights. assume that 30 people have heard the rumor as of today and that each day the number of people who have heard the rumor triples. the equation is 30(3)^t, predict when all americans will have heard the rumor. assume the population is 305 million

OpenStudy (anonymous):

365(30-305000)

OpenStudy (anonymous):

set 30(3^t) = 305 million. divide both sides by 30 to get 3^t = 305mil/30. take the natural log of both sides and use properties of exponents to bring the t in front of the natural log like this: t*ln(3) = ln(305mil/30). then divide both sides by ln(3) to get t

OpenStudy (anonymous):

i got 14.69 days

OpenStudy (anonymous):

ok let me see if i get it. hold on

OpenStudy (anonymous):

i keep on getting 2.1109

OpenStudy (anonymous):

i knew that :P

OpenStudy (anonymous):

i am doing something wrong???

OpenStudy (anonymous):

what do you get when you do 305 million / 30? it should be 10166666.67. then take the natural log of that. you should get 16.13. divide by ln(3) to get 14.68

OpenStudy (anonymous):

i got the answer 10166666.67but after that i dont get 14.68

OpenStudy (anonymous):

you need to take the natural log of 10166666.67 first

OpenStudy (anonymous):